OpenStudy (anonymous):

Simplify the expression 5x − (3x^2 + 9x).

OpenStudy (zehanz):

First, you have to write the expression without the brackets. To see what it will be, you can consider the "-" as "-1*": \(5x-(3x^2+9x)=5x-1\cdot(3x^2+9x)\). Now you can use the distributive property: a(b+c)=ab+ac to get rid of the brackets. So, multiply -1 with both the numbers 3x² and 9x.